High Desert Potteryworks makes a variety of pottery products that it sells to retailers. The company uses a job-order costingReq 1 Req 2 Req ЗА Req 3B Compute the total overhead cost applied to Job 205. (Round Predetermined overhead rate to 2 decimRequired: 1. Compute the predetermined overhead rates used in the Molding Department and the Painting Department. 2. ComputeComplete this question by entering your answers in the tabs below. Req 1 Req 2 Req ЗА Req 3B If the job contained 35 units, w
Answer #1

Answer 1. Predetermined overhead rate

Molding $5.20 per MH
Painting $13.50 per DLH

Explanation

Predetermined overhead rate for Molding =

(Estimated fixed manufacturing overhead cost / Estimated machine-hours) + Variable manufacturing overhead per machine-hour

= ($228,800 / 88,000 MH) + $2.60 = $5.20 per MH

Predetermined overhead rate for Molding =

(Estimated fixed manufacturing overhead cost / Estimated direct labor-hourss) + Variable manufacturing overhead per irect labor-hours

= ($445,890 / 50,100 MH) + $4.60 = $13.50 per MH

Answer 2.

Total overhead cost $3,720

Explanation

Molding Department overhead applied (370 MH * $5.20) 1,924
Painting Department overhead applied (133 DLH * $13.50) 1,796
Total overhead cost $3,720

Answer 3 A.

Molding Painting Total
Direct materials 940 1,160 $2,100
Direct labor 720 960 1.680
Manufacturing overhead 1,924 1,796 3,720
Total cost $3,584 $3,916 $7,500

Answer 3 B.

Unit product cost $214.29 per unit

Explanation

Unit product cost = Total manufacturing cost / Number of units

= $7,500 / 35 units = $214.29.